3x-2y=10, 2x+3y=-2 by substitution
Answer by richwmiller(17219)

substitution is the harder way to go. This was made for elimination or subtraction/addition
3x-2y=10
2x+3y=-2
First by elimination or addition
multiply first by 3 and second by 2
9x-6y=30
4x+6y=-4
add
13x=26
x=2
2x+3y=-2
4+3y=-2
3y=-6
y=-2
(2,-2)
now by substitution
3x-2y=10
2x+3y=-2
solve first for x
3x=10+2y
x=10/3+2y/3
plug it in the second
2*(10/3+2y/3)+3y=-2
20/3+4y/3+3y=-2
multiply by 3
20+4y+9y=-6
13y=-26
y=-2
plug it in any equation and solve for x
3x-2y=10
3x-2*(-2)=10
3x+4=10
3x=6
x=2